OpenStudy (anonymous):

MEDAL a bullet is fired horizontally at a velocity of 220 m/s. it strikes the ground at a distance of 511 m from where it was fired. how high above the ground was the bullet when it fired

OpenStudy (nethercreep333):

Whats the question?

OpenStudy (anonymous):

how high above the ground was the bullet when it fired

OpenStudy (anonymous):

physics

OpenStudy (amistre64):

start with maybe your linear equations of motion ....

OpenStudy (amistre64):

if its fired straight out, there is no initial velocity up/down and there is no acceleration acting on it from side to side since gravity pulls down, not out

OpenStudy (anonymous):

d=vt

OpenStudy (amistre64):

good, so t = d/v what was our time?

OpenStudy (anonymous):

2.3 s

OpenStudy (amistre64):

how far does something drop in 2.3 seconds?

OpenStudy (amistre64):

1/2 gt^2 right?

OpenStudy (anonymous):

is 2.3 correct

OpenStudy (amistre64):

its good enough for me, but im not the one grading it. take it to however many decimals you want to

OpenStudy (anonymous):

now do we use d=vit+1/2at^2

OpenStudy (amistre64):

or just let it ride at 511/220

OpenStudy (amistre64):

we can yes, vi=0 in this case

OpenStudy (amistre64):

it more appropriate as: vi sin(theta) t sin(0) = 0 as its shot horizontally

OpenStudy (anonymous):

use vi sin(theta) t for the formula

OpenStudy (amistre64):

that is a better generalization yes

OpenStudy (amistre64):

\[h(t)=\frac12gt^2+v_i~t~\sin(\theta) \]

OpenStudy (amistre64):

theta represent the angle of inclination from the horizontal

OpenStudy (anonymous):

so what do i plug in into the equation

OpenStudy (amistre64):

your values for gravity and time

OpenStudy (anonymous):

what is the intial velocity is it 0

OpenStudy (amistre64):

since we are shot horizontally, then the degrees above horizontal is 0 .. 211 * sin(0) is your initial velocity in the up/down directions

OpenStudy (anonymous):

so it is 0

OpenStudy (amistre64):

yeah :)

OpenStudy (amistre64):

so height = 1/2 gt^2

OpenStudy (anonymous):

h(t) = 1/2(-9.8)(2.3)^2+0(2.3)+sin(0)

OpenStudy (amistre64):

h(t) = 1/2(-9.8)(2.3)^2+211(2.3) sin(0)

OpenStudy (amistre64):

h(t) = 1/2(-9.8)(2.3)^2+0 h(t) = 1/2(-9.8)(2.3)^2

OpenStudy (anonymous):

why 211 should be 511

OpenStudy (amistre64):

hmm, should be 220 i spose but i mistyped it since i didnt feel like scrolling all the way up to verify that it was correctly inputed.

OpenStudy (amistre64):

but its immaterial, that term goes to 0

OpenStudy (amistre64):

the full generalization equation for distance an object falls is: \[h(t)=-\frac12gt^2+v_it~\sin(\theta )+h_i\] h(2.3) = 0, we are on the ground \[0=-\frac12g(2.3)^2+v_i(2.3)~\sin(0 )+h_i\] \[\frac12g(2.3)^2=h_i\]

OpenStudy (anonymous):

g=9.8

OpenStudy (amistre64):

yes, or whatever the gravity of the planet your testing it on happens to be :)

OpenStudy (anonymous):

25.92

OpenStudy (anonymous):

i mean -25.92

OpenStudy (amistre64):

and the dimensions is?

OpenStudy (anonymous):

horizontal

OpenStudy (amistre64):

no, initial height for this will be positive. you werent below ground to start with and it floated upwards ...

OpenStudy (amistre64):

dimensions refered to feet, meters, yards, kilometers, etc ...

OpenStudy (anonymous):

25.92 meters

OpenStudy (amistre64):

good

OpenStudy (anonymous):

is that the final answer

OpenStudy (amistre64):

what is it we were looking for?

OpenStudy (anonymous):

how high above the ground was the bullet when it fired

OpenStudy (amistre64):

and did we find that?

OpenStudy (anonymous):

i think so but not sure

OpenStudy (amistre64):

h = 1/2 gt^2 so yeah, we found the h

OpenStudy (anonymous):

can i ask another question

OpenStudy (amistre64):

sure

OpenStudy (anonymous):

A ball is thrown horizontally with an initial velocity of 20.0 meters per second from the top of a tower 60.0 high. What is the approximate total time required for the ball to reach the ground.

OpenStudy (amistre64):

what was our generalization of finding height?

OpenStudy (anonymous):

height equal 25.92

OpenStudy (amistre64):

no, that was a specific case, not the general formula that i posted.

OpenStudy (anonymous):

im on the next question

OpenStudy (amistre64):

the great thing about generalizations, is that they work regardless of the question that you are on .. you simply reuse them with the new specifics.

OpenStudy (anonymous):

d=vit+1/2at^2

OpenStudy (amistre64):

almost, but then vi might confuse people if they dont understand that it pertains only to the up/down direction ..... hence the sin(theta)

OpenStudy (anonymous):

so d=sin(theta) +1/2at^2

OpenStudy (amistre64):

no, we dont eliminate the rest of it in favor of sin(theta) the forces involved are this:  vi is important, and so is sin(theta). sin(theta) tells us how much of the initial velocity is imparted up/down